OpenStudy (korosh23):

A normal adult breathes in and exhales about 0.84 litres of air every 4 seconds with the minimal amount in the lungs of 0.08 Litres. Write a COSINE equation with 0≤t≤8 and find the time of maximum air capacity in this interval.

Miracrown (miracrown):

Suppose the total capacity of lungs be V0 Now, let y be the amount of air in lungs at any moment t Then, y(t)= a + (b* cos(w*t)) here a,b,w are unknown constants We need to determine their value from the given information; the minimum amount in lungs= 0.08 the minimum of the equation on board is a-b hence, we get to write, a-b = 0.08 Agree thus far? :)

OpenStudy (korosh23):

How do you know if a - b equals the minimum amount. a = amplitude b= horizontal compression or expansion of the graph.

Miracrown (miracrown):

a= constant amount b= amplitude when y obtains minimum value, cos(wt)= -1 so, y becomes a-b

OpenStudy (korosh23):

No, please write it in this form if it is possible. y= a cos(b (x -c )) +d a = amplitude b= horizontal expansion or compression c = phase shift d = vertical shift or translation It is easier for me to understand what you tell me.

OpenStudy (korosh23):

I agree to your minimum value = 0.08 L \[\left| \min \right| = d + a\]

OpenStudy (korosh23):

yes my bad 0.08 = d +a

OpenStudy (korosh23):

b = 2pi /4 or pi / 2

Miracrown (miracrown):

Right. The first thing to note is that there is no phase shift we are considering here, so c=0 now, min is 0.08 = d -a min is NOT d+a d+ a will be the max now, as I said earlier, it takes 4 seconds to repeat the cycle. hence T= 4 in our form, x stands for time t now, T=4= 2pi/b

Miracrown (miracrown):

so we get b= 2pi/4 = pi/2 A normal adult breathes in and exhales about 0.84 litres of air every 4 seconds b= pi/2 now, we need to write for 0.84 0.84 is the difference between the max and the min. The min value was already found to be d -a ... the max value will be d+a

Miracrown (miracrown):

the difference between max and min is (a+d) - (d-a) which turns out to be 2a and that equals 0.84 2a= 0.84 , agree?

OpenStudy (korosh23):

perfect :)

Miracrown (miracrown):

Alright, now solve for a,d we got 2a=0.84 we got d-a = 0.08 so, what are a, d?

OpenStudy (korosh23):

a= 0.42 d= 0.42 + 0.08 = 0.50

Miracrown (miracrown):

Right! :) and we already found c=0 b=pi/2 so, can you now write the equation?

OpenStudy (korosh23):

y = 0. 42 cos [ pi/2 (t)] + 0.50

Miracrown (miracrown):

Exactly, you got it! x'D

OpenStudy (korosh23):

find the time of maximum air capacity in this interval.

OpenStudy (korosh23):

Is it 0 and 4?

Miracrown (miracrown):

Oh... that happens when cos(bt)=1 So, t=0 is definitely a choice and t=4 is also a choice :)

Miracrown (miracrown):

since 0≤t≤8 t=8 is also a choice

Miracrown (miracrown):

so t=0,4,8

OpenStudy (korosh23):

but I have a very weird answer of it from my teacher which does not make sense I will write it in a sec

OpenStudy (korosh23):

Maximum capacity is when cos (pi . t /2)= -1 pi . t/2 = pi --> t =2 pi .t /2 = 3 pi --> t =6 Air capacity is maximum at 2 sec and 6 sec.

Miracrown (miracrown):

I do not understand that logic. I would think that t=2 and t=6 corresponds to minimum air capacity.

Miracrown (miracrown):

I think you should ask this to your teacher...

OpenStudy (korosh23):

OK, but we managed to go through most of it. I appreciate your help :)

Miracrown (miracrown):

:)

OpenStudy (perl):

*